# Fractional Deep Reinforcement Learning for Age-Minimal Mobile Edge Computing

### 16 Dec 2023 v1

@@ 1 Introduction
While the majority of existing studies on MEC were concerned about delay reduction (e.g., (Wang et al. 2021; Tang and Wong 2022)), most of real-time applications mentioned above concern about fresh status updates, while delay itself does not directly reflect timeliness. Here we highlight the huge difference between delay and AoI. Specifically, task delay takes into account only the duration between when the task is generated and when the task output has been received by the mobile device. Thus, under less frequent updates (i.e., when tasks are generated in a lower frequency), task delays are naturally smaller. This is because infrequent updates lead to empty queues and hence reduced queuing delays of the tasks. In contrast, AoI takes into account both the task delay and the freshness of the task output. Thus, to minimize the AoI with computational-intensive tasks, the update frequency needs to be neither too high nor too low in order to reduce the delay of each task while ensuring the freshness of the most up-to-date task output. [^1] Interesting readers may refer to (Talak and Modiano 2021)(Talak and Modiano 2021) for a more detailed analysis of tradeoffs between AoI and delay. More importantly, such a difference between delay and AoI leads to a counter-intuitive important phenomenon in designing age-minimal scheduling policy: upon the reception of each update, the mobile device may need to wait for a certain amount of time to generate the next new task (Sun et al. 2017).
@@ 2 Literature Review
Age of Information: Kaul et al. first introduced AoI in (Kaul, Yates, and Gruteser 2012). Assuming complete and known statistical information, the majority of this line of work mainly focused on the optimization and analysis of AoI in queueing systems and wireless networks (see (Zou, Ozel, and Subramaniam 2021; Chiariotti et al. 2021; Kuang et al. 2020), and a survey in (Yates et al. 2021)). Zou et al. in (Zou, Ozel, and Subramaniam 2021), Chiariotti et al. in (Chiariotti et al. 2021), and Kuang et al. in (Kuang et al. 2020). The above studies analyzed simple single-device-single-server models and hence did not consider offloading.
@@ 3 System Model
Scheduler: At the time when task $k$ of mobile device $m\in\mathcal{M}$ is generated (i.e., time $t_{m,k}$), the task scheduler observes the queue lengths of edge nodes and makes the offloading decision denoted by $x_{m,k}\in\{0\}\cup\mathcal{N}$. Let $\bm{s}^{\textsc{o}}_{m}(k)$ denote the state vector associated with task $k$ of mobile device $m$: [^2] Under the system model in Section 3, observing the queue lengths of edge nodes is sufficient for mobile devices to learn their offloading policies. Under a more complicated system (e.g., with device mobility), mobile devices may need to observe additional state information, while our proposed fractional RL framework and DRL algorithm are still applicable with the extended state vector. $\bm{s}^{\textsc{o}}_{m}(k)=\bm{q}(t_{m,k}),~{}k\in\mathcal{K},m\in\mathcal{M}$, where $\bm{q}(t_{m,k})=(q_{n}(t_{m,k}),n\in\mathcal{N})$ corresponds to the queue lengths of all edge nodes. We assume that edge nodes send their queue length information upon the requests of mobile devices. Since a generator generates a new task only after the previous task has been processed, the queue length is less than or equal to $M$. Thus, it can be encoded in $O(\log_{2}M)$ bits, which incurs only small signaling overheads. Let $\mathcal{S}^{\textsc{o}}=\mathcal{M}^{1\times N}$ denote the state space. Let $\bm{a}^{\textsc{o}}_{m}(k)$ denote the action associated with task $k$ of mobile device $m$. Thus, $\bm{a}^{\textsc{o}}_{m}(k)=x_{m,k},~{}~{}k\in\mathcal{K},m\in\mathcal{M}$. Let $\mathcal{A}^{\textsc{o}}\in\{0\}\cup\mathcal{N}$ denote the offloading action space. Let ${\pi}^{\textsc{o}}_{m}$ denote the task offloading policy of mobile device $m\in\mathcal{M}$.
@@
We aim at the optimal policy $\bm{\pi}_{m}^{*}$ for each mobile device $m\in\mathcal{M}$ to minimize its expected time-average AoI, i.e., [^3] If the AoI minimization problem is modeled as a game or multi-agent RL-based approaches (e.g., (Chen et al. 2022)), the policies of mobile devices will converge to Nash equilibrium rather tan reduce the AoI of each device, and they need additional signaling between mobile devices for state information sharing. Thus, in this work, to minimize time-average AoI, we model the policy design problem of each mobile device as an optimization problem (rather than game). The simulation results in Section 6 show that with our proposed approach, the scheduling policies of mobile devices can gradually converge.
@@ 4 Fractional RL Framework
Therefore, the discounted reformulation in (10) serves as a good approximation of (7) when $\delta$ approaches $1$. The main purpose of introducing the discounted reformulation in (10) is that the majority of existing studies on RL (and DRL) algorithm design have focused on the discounted formulation (e.g., (Tang and Wong 2022; Zhao et al. 2022; Nguyen et al. 2021; Chen et al. 2022)). Hence, such an approximation provides us with a convention of designing new DRL algorithms for fractional MDP problems based on existing well-established DRL algorithms. We will therefore stick to the discounted reformulation for the rest of this paper.
@@
A fractional MDP is defined as $(\mathcal{S},\mathcal{A},P,c_{N},c_{D},\delta)$, where $\mathcal{S}$ and $\mathcal{A}$ are the finite sets of states and actions, respectively; $P$ is the transition distribution; $c_{N}$ and $c_{D}$ are the cost functions, [^4] Since we aim at minimizing the time-average AoI, we consider minimizing long-term expected cost in this work. The proposed approach can be easily extended to the fraction MDP that maximizes long-term expected reward. and $\delta$ is a discount factor. We use $\mathcal{Z}$ to denote the joint state-action space, i.e., $\mathcal{Z}\triangleq\mathcal{S}\times\mathcal{A}$.
@@
where we can see from Lemmas 1 and 2 that $\gamma^{*}$ satisfies [^5] Note that $\gamma^{*}$ always exists due to the minimization nature.
@@
While the convergence proof in (Dinkelbach 1967) requires to obtain the exact solution in each episode, Theorem 1 generalizes this result to the case where we only obtain an approximated (inexact) solution in each episode. In addition to the proof techniques in (Dinkelbach 1967) and (Ghavamzadeh et al. 2011), our proof techniques include induction and exploiting the convexity of $Q_{i}(\bm{s},\bm{a})$. We present a proof sketch of Theorem 1 in Appendix A.
@@
Although as $\{Q_{i}(\bm{s}_{0},\bm{a}_{i})\}$ is convergent to 00 and hence $\epsilon_{i}<-\alpha Q_{i}$ is getting more restrictive as $i$ increases, the steps needed $T_{i}$ in Algorithm 1 keep to be finite without increasing over episode $i$ based on (Ghavamzadeh et al. 2011). See Appendix B in detail.
@@
5 Fractional DRL Algorithm
In this section, we present a fractional DRL-based algorithm to approximate the Q-function in FQL algorithm and solve Problem (6) with DDPG (for continuous action space) (Lillicrap et al. 2015) and D3QN (for discrete action space) (Mnih et al. 2015) techniques [^6] Twin Delayed DDPG (TD3) (Fujimoto, van Hoof, and Meger 2018) and Proximal policy optimization (PPO) (Schulman et al. 2017) are also applicable to continuous action. Soft Actor-Critic for Discrete Action (SAC-discrete) (Christodoulou 2019) also works for discrete action space. We empirically show via extensive experiments that in our algorithm outperforms the other aforementioned techniques. Note that our proposed algorithm can be extended by incorporating other techniques. for the task updating and offloading processes in the decentralized manner, which is illustrated in Fig. 2. Appendix C shows detailed settings of networks with the reference of settings in (Tang and Wong 2022, Table I). Moreover, we design a cost function based on our fractional RL framework to ensure the convergence.
@@ 5 Fractional DRL Algorithm
where $Y_{m,l}^{i}$ and $Z_{m,l+1}^{i}$ are the delay of task $l$ and the time that the task generator waits for generating the next task $l+1$, respectively, for mobile device $m$. Note that $Y_{m,l}^{i}$ is a function of $\bm{a}^{\textsc{h}}_{m,i}(l)$, and $Z_{m,l+1}^{i}$ is a function of $\bm{a}^{\textsc{h}}_{m,i}(l)$ and $\bm{a}^{\textsc{u}}_{m,i}(l+1)$. The cost module keeps track of $\bm{a}^{\textsc{h}}_{m,i}(k)$ or equivalently $Y^{i}_{m,l}$ and $Z^{i}_{m,l}$ for all $l=1,...,k$ across the training process.
In step $k$ of episode $i$, [^7] Here, step $k$ refers to the updating and offloading processes as well as the training process associated with task $k$. a cost is determined and sent to the DDPG and D3QN modules. This process corresponds to the inner loop of the proposed fractional RL framework and is defined based on (21): for all $i\in[E],m\in\mathcal{M},k\in\mathcal{K},$
@@ 6 Performance Evaluation
We denote our proposed approach with Frac. DRL (OFL+U), which is short for “fractional DRL with offloading and updating policies”. This is compared with several benchmark methods:
@@
Edge Capacity: Fig. 4(a) evaluates the performance of our proposed schemes under different node processing capacities. First, the proposed fractional DRL-based algorithm consistently achieves lower average AoI, compared against those non-fractional benchmarks. Such an advantage is more significant as the per-node processing capacity is larger. When processing capacity is $75$ GHz, Frac. DRL (OFL+U) can achieve an average AoI reduction of $54.8\%$ and $73.6\%$, compared against PGOA and Non-Frac. DRL, respectively. Second, when compared with Frac. DRL (OFL), Frac. DRL (OFL+U) can further reduce the average AoI up to $19.9\%$ when the processing capacity of edge nodes is 55 GHz. This further shows that, in addition to the fractional RL framework, a well-designed updating policy also plays an important role of further improving the performance, especially there are relatively high edge loads.
Drop Coefficient: In Fig. 4(b), we consider different drop coefficients, i.e., the ratio of the drop time $\bar{Y}$ to the average time of processing a task. The performance gaps between the proposed schemes and benchmarks are large when the drop ratio gets smal (i.e., tasks are more delay-sensitive). When the drop coefficient is 1.6, the Frac. DRL (OFL+U) reduces the average AoI by $55.1\%$, compared with Non-Frac. DRL.
@@
7 Conclusion
This paper has studied the computational task scheduling (including offloading and updating) problem for age-minimal MEC. To address the underlying challenges of unknown load dynamics and the fractional objective, we have proposed a fractional RL framework with a provable linear convergence rate. We further designed a fractional DRL algorithm that incorporates D3QN and DDPG techniques to tackle hybrid action space. Experimental results show that our proposed fractional algorithms significantly reduce the average AoI, compared against several benchmarks. Meanwhile, the joint optimization of offloading and updating can further reduce the average AoI, validating the effectiveness of our proposed scheme. There are several future directions, including incorporating RL techniques for non-stationary environments and social optimal scheduling.
@@ 7 Conclusion
9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 19 Dec 2023 v2

@@ Introduction
While the majority of existing studies on MEC were concerned about delay reduction (e.g., (Wang et al. 2021; Tang and Wong 2022)), most of real-time applications mentioned above concern about fresh status updates, while delay itself does not directly reflect timeliness. Here we highlight the huge difference between delay and AoI. Specifically, task delay takes into account only the duration between when the task is generated and when the task output has been received by the mobile device. Thus, under less frequent updates (i.e., when tasks are generated in a lower frequency), task delays are naturally smaller. This is because infrequent updates lead to empty queues and hence reduced queuing delays of the tasks. In contrast, AoI takes into account both the task delay and the freshness of the task output. Thus, to minimize the AoI with computational-intensive tasks, the update frequency needs to be neither too high nor too low in order to reduce the delay of each task while ensuring the freshness of the most up-to-date task output. More importantly, such a difference between delay and AoI leads to a counter-intuitive important phenomenon in designing age-minimal scheduling policy: upon the reception of each update, the mobile device may need to wait for a certain amount of time to generate the next new task (Sun et al. 2017).
@@
What’s more, the age-minimal technique can also be extended to other ratio optimization problems including financial portfolio optimization (Keating and Shadwick 2002) and energy efficiency (EE) maximization problems for wireless communications (Omidkar et al. 2022).
@@
Age of Information: Kaul et al. first introduced AoI in (Kaul, Yates, and Gruteser 2012). Assuming complete and known statistical information, the majority of this line of work mainly focused on the optimization and analysis of AoI in queueing systems and wireless networks (see (Zou, Ozel, and Subramaniam 2021; Chiariotti et al. 2021; Kuang et al. 2020), and a survey in (Yates et al. 2021)). Zou et al. in (Zou, Ozel, and Subramaniam 2021), Zhou et al. in (M. Zhou and Yates 2024, Early Access), Chiariotti et al. in (Chiariotti et al. 2021), and Kuang et al. in (Kuang et al. 2020). The above studies analyzed simple single-device-single-server models and hence did not consider offloading.
@@
Scheduler: At the time when task $k$ of mobile device $m\in\mathcal{M}$ is generated (i.e., time $t_{m,k}$), the task scheduler observes the queue lengths of edge nodes and makes the offloading decision denoted by $x_{m,k}\in\{0\}\cup\mathcal{N}$. Let $\bm{s}^{\textsc{o}}_{m}(k)$ denote the state vector associated with task $k$ of mobile device $m$: $\bm{s}^{\textsc{o}}_{m}(k)=\bm{q}(t_{m,k}),~{}k\in\mathcal{K},m\in\mathcal{M}$, where $\bm{q}(t_{m,k})=(q_{n}(t_{m,k}),n\in\mathcal{N})$ corresponds to the queue lengths of all edge nodes. We assume that edge nodes send their queue length information upon the requests of mobile devices. Since a generator generates a new task only after the previous task has been processed, the queue length is less than or equal to $M$. Thus, it can be encoded in $O(\log_{2}M)$ bits, which incurs only small signaling overheads. Let $\mathcal{S}^{\textsc{o}}=\mathcal{M}^{1\times N}$ denote the state space. Let $\bm{a}^{\textsc{o}}_{m}(k)$ denote the action associated with task $k$ of mobile device $m$. Thus, $\bm{a}^{\textsc{o}}_{m}(k)=x_{m,k},~{}~{}k\in\mathcal{K},m\in\mathcal{M}$. Let $\mathcal{A}^{\textsc{o}}\in\{0\}\cup\mathcal{N}$ denote the offloading action space. Let ${\pi}^{\textsc{o}}_{m}$ denote the task offloading policy of mobile device $m\in\mathcal{M}$.
@@
We aim at the optimal policy $\bm{\pi}_{m}^{*}$ for each mobile device $m\in\mathcal{M}$ to minimize its expected time-average AoI.
@@
Therefore, the discounted reformulation in (10) serves as a good approximation of (7) when $\delta$ approaches $1$. Such an approximation provides us with a convention of designing new DRL algorithms for fractional MDP problems based on existing well-established DRL algorithms. We will stick to the discounted reformulation for the rest of this paper.
@@
A fractional MDP is defined as $(\mathcal{S},\mathcal{A},P,c_{N},c_{D},\delta)$, where $\mathcal{S}$ and $\mathcal{A}$ are the finite sets of states and actions, respectively; $P$ is the transition distribution; $c_{N}$ and $c_{D}$ are the cost functions ,  [^1] Since we aim at minimizing the time-average AoI, we consider minimizing long-term expected cost in this work. and $\delta$ is a discount factor. We use $\mathcal{Z}$ to denote the joint state-action space, i.e., $\mathcal{Z}\triangleq\mathcal{S}\times\mathcal{A}$.
@@
where we can see from Lemmas 1 and 2 that $\gamma^{*}$ satisfies
@@
While the convergence proof in (Dinkelbach 1967) requires to obtain the exact solution in each episode, Theorem 1 generalizes this result to the case where we only obtain an approximated (inexact) solution in each episode. In addition to the proof techniques in (Dinkelbach 1967) and (Ghavamzadeh et al. 2011), our proof techniques include induction and exploiting the convexity of $Q_{i}(\bm{s},\bm{a})$. We present a proof sketch of Theorem 1 in Appendix A [^2] Please refer to (Jin et al. 2023) for our appendices..
@@
Although as $\{Q_{i}(\bm{s}_{0},\bm{a}_{i})\}$ is convergent to 00 and hence $\epsilon_{i}<-\alpha Q_{i}$ is getting more restrictive as $i$ increases, the steps needed $T_{i}$ in Algorithm 1 keep to be finite without increasing over episode $i$. See Appendix B in detail.
@@
Fractional DRL Algorithm
In this section, we present a fractional DRL-based algorithm to approximate the Q-function in FQL algorithm and solve Problem (6) with DDPG (for continuous action space) (Lillicrap et al. 2015) and D3QN (for discrete action space) (Mnih et al. 2015) techniques for the task updating and offloading processes in the decentralized manner, which is illustrated in Fig. 2. Appendix C shows detailed settings of networks. Moreover, we design a cost function based on our fractional RL framework to ensure the convergence.
@@
where $Y_{m,l}^{i}$ and $Z_{m,l+1}^{i}$ are the delay of task $l$ and the wait interval for generating the next task $l+1$, respectively, for mobile device $m$. Note that $Y_{m,l}^{i}$ is a function of $\bm{a}^{\textsc{h}}_{m,i}(l)$, and $Z_{m,l+1}^{i}$ is a function of $\bm{a}^{\textsc{h}}_{m,i}(l)$ and $\bm{a}^{\textsc{u}}_{m,i}(l+1)$. The cost module keeps track of $\bm{a}^{\textsc{h}}_{m,i}(k)$ or equivalently $Y^{i}_{m,l}$ and $Z^{i}_{m,l}$ for all $l=1,...,k$ across the training process.
In step $k$ of episode $i$, a cost is determined and sent to the DDPG and D3QN modules. This process corresponds to the inner loop of the proposed fractional RL framework and is defined based on (21): for all $i\in[E],m\in\mathcal{M},k\in\mathcal{K},$
@@
We denote proposed Frac. DRL (OFL+U), which is short for “fractional DRL with offloading and updating policies”. This is compared with several benchmark methods:
@@
Edge Capacity: Fig. 4(a) evaluates the performance of our proposed schemes under different node processing capacities. First, the proposed fractional DRL-based algorithm consistently achieves lower average AoI, compared against those non-fractional benchmarks. Such an advantage is more significant as the per-node processing capacity is larger. When processing capacity is $75$ GHz, Frac. DRL (OFL+U) can achieve an average AoI reduction of $54.8\%$ and $73.6\%$, compared against PGOA and Non-Frac. DRL, respectively. Second, when compared with Frac. DRL (OFL), Frac. DRL (OFL+U) can further reduce the average AoI up to $19.9\%$ when the processing capacity of edge nodes is 55 GHz. This further shows that a well-designed updating policy also plays an important role of further improving the performance, especially there are relatively high edge loads.
Drop Coefficient: In Fig. 4(b), we consider different drop coefficients, i.e., the ratio of the drop time $\bar{Y}$ to the average time of processing a task. The performance gaps between the proposed schemes and benchmarks are large when the drop ratio gets smal (i.e., tasks are more delay-sensitive). When the drop coefficient is 1.0, the Frac. DRL (OFL+U) reduces the average AoI by $57.6\%$, compared with Non-Frac. DRL.
@@
Conclusion
This paper has studied the computational task scheduling (including offloading and updating) problem for age-minimal MEC. To address the underlying challenges of unknown load dynamics and the fractional objective, we have proposed a fractional RL framework with a provable linear convergence rate. We further designed a fractional DRL algorithm that incorporates D3QN and DDPG techniques to tackle hybrid action space. Experimental results show that proposed fractional algorithms significantly reduce the average AoI, compared against several benchmarks. Meanwhile, the joint optimization of offloading and updating can further reduce the average AoI, validating the effectiveness of our proposed scheme. There are several future directions, including incorporating multi-agent RL with recurrent neural networks for non-stationarity and social optimal scheduling.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Projects 62202427 and 62202214), in part by Guangdong Basic and Applied Basic Research Foundation under Grant 2023A1515012819, and in part by the Australian Research Council (ARC) Discovery Early Career Researcher Award (DECRA) under Grant DE230100046.